a) 0.056m³ of air at 1.38 bar are compressed isothermally to a volume of 0.014m³. 

Determine the work energy required for the compression.


b) compare this to the work energy which would be required for an adibatic compression through the same volume ratio. For air R= 0.287KN and ➰ = 1.4


1
Expert's answer
2022-02-18T08:22:18-0500

Given:

"V_1=0.056\:\rm m^3"

"V_2=0.014\:\rm m^3"

"p_1=1.38*10^5\:\rm Pa"


(a) "T=const"

"W=p_1V_1\ln\frac{V_2}{V_1}"

"W=1.38*10^5*0.056\ln\frac{0.014}{0.056}=-10713\:\rm J"

(b) "\delta Q=0"

"W=\frac{p_1V_1((\frac{V_2}{V_1})^{1-\gamma}-1)}{1-\gamma}"

"W=\frac{1.38*10^5*0.056((\frac{0.014}{0.056})^{1-1.4}-1)}{1-1.4}"

"W=-14318\:\rm J"